A 5.50-g sample of CaO is reacted with 5.31 g of H2O. How many grams of water remain after the reaction is complete?
1
Expert's answer
2020-12-21T03:38:34-0500

CaO + H2O = Ca(OH)2

n(CaO) = m/M = 5.5g/56g/mol = 0.098 mol

n(H2O) = 5.31g/18g/mol = 0.295 mol

According to the equation

n(CaO) : n(H2O) = 1:1

CaO reacts completely, H2O is in the excess.

0.098 mol H2O reacts with CaO.

n(H2O) = 0.295 - 0.098 = 0.197 mol of water will remain after the reaction is complete.

m(H2O) = 0.197mol*18g/mol = 3.546 g